# µGUIDE: a framework for microstructure imaging via generalized uncertainty-driven inference using deep learning

### v1

@@ 2.1.1 The inference problem
Forward models are designed to mimic at best a given biophysical phenomenon, for some given time and scale (Alexander, b; Yablonskiy and Sukstanskii, ; Jelescu and Budde, ; Novikov et al., a; Alexander et al., ; Jelescu et al., b). As a consequence, forward models are injection functions (every biologically plausible $\boldsymbol{\theta_{i}}$ generates exactly one signal $\boldsymbol{x_{i}}$), but do not always happen to be bijections, meaning that multiple $\boldsymbol{\theta_{i}}$ can generate the same signal $\boldsymbol{x_{i}}$. It can be impossible, based on biological considerations, to infer which solution $\boldsymbol{\theta_{i}}$ best reflects the probed structure. We refer to these models as ’degenerate models’.
Point estimates algorithms, such as minimum least square or maximum likelihood estimation algorithms, allow to estimate one set of microstructure parameters that could explain an observed signal. In the case of degenerate models, the solution space can be multi-modal and those algorithms will hide possible solutions. When considering real-life acquisitions, i.e. noisy and/or under-sampled acquisitions, one also needs to consider the bias introduced with respect to the forward model, and the resulting variance in the estimates (Jones, ; Behrens et al., ).
@@
An analytical expression of the likelihood distribution is often hard to write for data points generated by complex non-linear models, making difficult the use of Equation 2. Models that do not admit a tractable likelihood are called implicit models (Diggle and Gratton, ). Two approaches have then been developed to allow for the estimation of this posterior distribution: either by approximating the likelihood distribution, or by learning directly the posterior distribution using a conditional density estimator. While this work focuses on the estimate of the posterior distribution using a conditional density estimator, we show a comparison with MCMC, which are commonly used methods in the community. We will therefore introduce this method in the following paragraph.
2.1.3 Estimating the likelihood function.
Well-established approaches for estimating the likelihood function are MCMC methods. These methods rely on a noise model to define the likelihood distribution, such as the Rician (Panagiotaki et al., a) or Offset Gaussian models (Alexander, b). In this work, we will be using the Microstructure Diffusion Toolbox to perform the MCMC computations (Harms and Roebroeck, ), which relies on the Offset Gaussian model. The log-likelihood function is then the following:
@@
where $\mathcal{M}(\boldsymbol{\theta})$ is the signal obtained using the biophysical model, $\sigma$ is the standard deviation of the Gaussian distributed noise, estimated from the reconstructed magnitude images (Dietrich et al., ), and $m$ is the number of observations in the dataset.
MCMC methods allow to obtain posterior distributions using Bayes’ formula (Eq. 2) with the previously defined likelihood function (Eq. 3) and some prior distributions, which are usually uniform distributions defined on biologically plausible ranges. They generate a multi-dimensional chain of samples which is guaranteed to converge towards a stationary distribution, which approximates the posterior distribution (Metropolis et al., ; Hastings, ).
The need to compute the signal following the forward model at each iteration makes these sampling methods computationally expensive and time consuming. In addition, they require some adjustments specific to each model, such as the choice of burn-in length, thinning and the number of samples to store. Harms and Roebroeck recommend to use the Adaptive Metropolis-Within-Gibbs (AMWG) algorithm for sampling dMRI models, initialized with a maximum likelihood estimator obtained from non-linear optimization, with 100 to 200 samples for burn-in and no thinning. The recommended number of samples is model-dependent. Authors recommendations can be found in their paper.
2.1.4 Bypassing the likelihood function.
An alternative method was proposed to overcome the challenges associated with approximating the likelihood function and the limitations of MCMC sampling algorithms. This approach involves directly approximating the posterior distribution by utilizing a conditional density estimator, i.e. a family of conditional probability density function approximators denoted as $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$. These approximators are parameterized by $\phi$ and accept both the parameters $\boldsymbol{\theta}$ and the observation $\boldsymbol{x}$ as input arguments. Our posterior approximation is then obtained by minimizing its average Kullback-Leibler divergence with respect to the conditional density estimator for different choices of $\boldsymbol{x}$, as per (Papamakarios and Murray, ):
@@
If the class of conditional density estimators is sufficiently expressive, it can be demonstrated that the minimizer of Eq. (6) converges to $p(\boldsymbol{\theta}|\boldsymbol{y})$ when $N\to\infty$ (Greenberg et al., ). It is worth noting that the parametrization $\phi$, obtained at the end of the optimization procedure, serves as an amortized posterior for various choices of $\boldsymbol{x}$. Hence, for a particular observation $\boldsymbol{x}_{0}$, we can simply use $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x}_{0})$ as an approximation of $p(\boldsymbol{\theta}|\boldsymbol{x}_{0})$.
3 Methods
3.1 Neural density estimators.
@@ 3 Methods
In this study, the Sequential Neural Posterior Estimation (SNPE-C) algorithm (Papamakarios and Murray, ; Greenberg et al., ) is employed to train a neural network that directly approximates the posterior distribution. Thus, sampling from the posterior can be done by sampling from the trained neural network. Neural density estimators have the advantage of providing exact density evaluations, in contrast to Variational Autoencoders (VAE (Kingma and Welling, )) or generative adversarial networks (GAN (Goodfellow et al., )), which are better suited for generating synthetic data.
@@
The conditional probability density function approximators used in this project belong to a class of neural networks called normalizing flows (Papamakarios et al., a). These flows are invertible functions capable of transforming vectors generated from a simple base distribution (e.g. the standard multivariate Gaussian distribution) into an approximation of the true posterior distribution. An autoregressive architecture for normalizing flows is employed, implemented via the Masked Autoregressive Flow (MAF (Papamakarios et al., b)), which is constructed by stacking five Masked Autoencoder for Distribution Estimation models (MADE (Germain et al., )).
@@
3.2 Handling the large dimensionality of the data
@@
As the dimensionality of the input data $\boldsymbol{x}$ grows, the complexity of the corresponding inverse problem also increases. Accurately characterizing the posterior distributions or estimating the tissue microstructure parameters becomes more challenging. As a consequence, it is often necessary to rely on a set of low-dimensional features (or summary statistics) instead of the raw data for the inference task process (Blum et al., ; Fearnhead and Prangle, ; Sisson et al., ; Papamakarios et al., c; Chen et al., ). Learning a set of sufficient statistics before estimating the posterior distribution makes the inference easier and offers many benefits (see e.g. the Rao-Blackwell theorem).
@@
A follow-up challenge lies in the choice of suitable summary statistics. For well-understood problems and data, it is possible to manually design these features using deterministic functions that condense the information contained in the raw signal into a set of handful summary statistics. Previous works, such as (Novikov et al., b) and (Jallais et al., ), have proposed specific summary statistics for two different biophysical models. However, defining these summary statistics is difficult and often requires prior knowledge of the problem at hand. In the context of dMRI, they also rely on acquisition constraints and are model-specific.
@@
In this work, the proposed framework aims to be applicable to any forward model and be as general as possible. We therefore propose to learn the summary statistics from the high-dimensional input signals $\boldsymbol{x}$ using a neural network. This neural network is referred to as an embedding neural network. The observed signals are fed into the embedding neural network, whose outputs are then passed to the neural density estimator. The parameters of the embedding network are learned together with the parameters of the neural density estimator, leading to the extraction of optimal features that minimize the uncertainty of $p(\boldsymbol{\theta}|\boldsymbol{x})$. Here, we propose to use a Multi-Layer Perceptron (MLP) with three layers as a summary statistics extractor. The number of features $N_{f}$ extracted by the MLP can be either defined a priori or determined empirically during training.
The full architecture of the proposed Bayesian framework, dubbed µGUIDE, is presented in Figure 1.
Figure 1: µGUIDE architecture. Input: an observed signal. Output: a posterior distribution of the model parameters. Based on a SBI framework, it combines a Multi-Layer Perceptron (MLP) with 3 layers and a normalizing flow. The MLP learns a low-dimensional representation of $\boldsymbol{x}$, based on a small number of features ($N_{f}$), that can be either defined a priori or determined empirically during training. The MLP is trained simultaneously with the normalizing flow, leading to the extraction of the optimal features that minimize the uncertainty of $p(\boldsymbol{\theta}|\boldsymbol{x})$, i.e. uncertainty-driven.
3.3 Quantifying the confidence in the estimates
To easily visualize the estimated posterior distributions in each voxel, it is necessary to define measures that quantify the best estimates and the associated confidence levels. These measures produce easily interpretable maps for each tissue microstructure parameter. We therefore propose the definition of three quantitative measures, extracted from the estimated posterior distributions.
We start by checking whether a posterior distribution is degenerate, that is if the distribution presents multiple distinct parameter solutions, appearing as multiple local maxima. In such cases, we highlight the corresponding voxel with a red dot on the estimated maps. For non-degenerate posterior distributions, we extract three quantities:
@@ 3.3 Quantifying the confidence in the estimates
Figure 2 presents those measures on exemplar posterior distributions.
Figure 2: Presentation of the measures introduced to quantify a posterior distribution on exemplar posterior distributions. Maximum A Posteriori (MAP): is the most likely parameter estimate (dashed vertical lines). Degeneracy: highlights the presence of multiple possible parameter solutions in a voxel. Uncertainty: measures the dispersion of the 50% most probable samples using the interquartile range, with respect to the prior range. Ambiguity: measures the Full Width at Half Maximum (FWHM), in percentage with respect to the prior range.
3.4 Biophysical Models
We show exemplar applications to three biophysical models from the literature with increased complexity. For each model, we compare the fitting quality of the posterior distributions obtained using the MLP and manually defined summary statistics.
  • Model 1: the Ball&Stick model (Behrens et al., ): This is a two compartment model. The brain tissue is modeled as a stick with diffusivity $D_{in}$ and signal fraction $f_{in}$, and a ball, with diffusivity $D_{e}$. The direction of the stick is randomly sampled on a sphere. This model has the main advantage of being non-degenerate. We define the summary statistics as the direction-averaged signal (seven b-shells, see section 3.6).
  • Model 2: the Standard Model (SM) (Novikov et al., a): The microstructure parameters of this two-compartment model are the neurite signal fraction $f$, the intra-neurite diffusivity $D_{a}$, the orientation dispersion index $ODI$, and the parallel and perpendicular diffusivities within the extra-neurite space $D_{e}^{\parallel}$ and $D_{e}^{\perp}$. We use the LEMONADE (Novikov et al., b) system of equations, which is based on a cumulant decomposition of the signal, to define six summary statistics.
  • Model 3: an extended-SANDI model (Palombo et al., ): The parameters of interest of this three-compartment model are the neurite signal fraction $f_{n}$, the intra-neurite diffusivity $D_{n}$, the orientation dispersion index $ODI$, soma signal fraction $f_{s}$, a proxy of soma radius and diffusivity $C_{s}$ (Jallais et al., ), extra-cellular signal fraction $f_{e}$ and isotropic diffusivity $D_{e}$. We use the six summary statistics defined in (Jallais et al., ), which are based on a high and low b-value signal expansion.

### v3

@@ 4.1.1 The inference problem
Forward models are designed to mimic at best a given biophysical phenomenon, for some given time and scale (Alexander, 2009; Yablonskiy and Sukstanskii, 2010; Jelescu and Budde, 2017; Novikov et al., 2018a; Alexander et al., 2019; Jelescu et al., 2020). As a consequence, forward models are injection functions (every biologically plausible $\boldsymbol{\theta_{i}}$ generates exactly one signal $\boldsymbol{x_{i}}$), but do not always happen to be bijections, meaning that multiple $\boldsymbol{\theta_{i}}$ can generate the same signal $\boldsymbol{x_{i}}$. It can be impossible, based on biological considerations, to infer which solution $\boldsymbol{\theta_{i}}$ best reflects the probed structure. We refer to these models as ’degenerate models’.
Point estimates algorithms, such as minimum least square or maximum likelihood estimation algorithms, allow to estimate one set of microstructure parameters that could explain an observed signal. In the case of degenerate models, the solution space can be multi-modal and those algorithms will hide possible solutions. When considering real-life acquisitions, i.e. noisy and/or under-sampled acquisitions, one also needs to consider the bias introduced with respect to the forward model, and the resulting variance in the estimates (Jones, 2003; Behrens et al., 2003).
@@ 4.1.2 The Bayesian formalism
To approximate these posterior distributions, traditional methods rely on the estimation of the likelihood $p(\boldsymbol{x_{0}}|\boldsymbol{\theta})$ of the observed data point $\boldsymbol{x_{0}}$ via an analytic expression. This likelihood function corresponds to an integral over all possible trajectories through the latent space, that is $p(\boldsymbol{x_{0}}|\boldsymbol{\theta})=\int p(\boldsymbol{x_{0}},\boldsymbol{z}|\boldsymbol{\theta})\mathrm{d}\boldsymbol{z}$, where $p(\boldsymbol{x_{0}},\boldsymbol{z}|\boldsymbol{\theta})$ is the joint probability density of observed data $\boldsymbol{x_{0}}$ and latent variables $\boldsymbol{z}$. For forward models with large latent spaces, computing this integral explicitly becomes impractical. The likelihood function is then intractable, rendering these methods unusable (Cranmer et al., 2020). Models that do not admit a tractable likelihood are called implicit models (Diggle and Gratton, 1984).
To circumvent this issue, some techniques have been proposed to sample numerically from the likelihood function, such as Markov-Chain-Monte-Carlo (MCMC) (Metropolis et al., 1953). Another set of approaches proposes to train a conditional density estimator to learn a surrogate of the likelihood distribution (Papamakarios et al., 2019; Lueckmann et al., 2019), the likelihood-ratio (Cranmer et al., 2016; Gutmann et al., 2018) or the posterior distribution (Papamakarios and Murray, 2016; Lueckmann et al., 2017; Papamakarios et al., 2019), allowing to greatly reduce computation times. These methods are dubbed Likelihood-Free Inference (LFI), or Simulation-Based Inference (SBI) methods (Cranmer et al., 2020; Tejero-Cantero et al., 2020). In particular, there has been a growing interest towards deep generative modeling approaches in the machine learning community (Lueckmann et al., 2021). They rely on specially tailored neural network architectures to approximate probability density functions from a set of examples. Normalizing flows (Papamakarios et al., 2021) are a particular class of such neural networks that have demonstrated promising results for SBI in different research fields (Gonçalves et al., 2020; Greenberg et al., 2019).
While this work focuses on the estimate of the posterior distribution using a conditional density estimator, we show a comparison with MCMC, which are commonly used methods in the community. We will therefore introduce this method in the following paragraph.
4.1.3 Estimating the likelihood function
Well-established approaches for estimating the likelihood function are MCMC methods. These methods rely on a noise model to define the likelihood distribution, such as the Rician (Panagiotaki et al., 2012) or Offset Gaussian models (Alexander, 2009). In this work, we will be using the Microstructure Diffusion Toolbox to perform the MCMC computations (Harms and Roebroeck, 2018), which relies on the Offset Gaussian model. The log-likelihood function is then the following:
@@ 4.1.3 Estimating the likelihood function
where $\mathcal{M}(\boldsymbol{\theta})$ is the signal obtained using the biophysical model, $\sigma$ is the standard deviation of the Gaussian distributed noise, estimated from the reconstructed magnitude images (Dietrich et al., 2007), and $m$ is the number of observations in the dataset.
MCMC methods allow to obtain posterior distributions using Bayes’ formula (Eq. (2)) with the previously defined likelihood function (Eq. (3)) and some prior distributions, which are usually uniform distributions defined on biologically plausible ranges. They generate a multi-dimensional chain of samples which is guaranteed to converge towards a stationary distribution, which approximates the posterior distribution (Metropolis et al., 1953).
The need to compute the signal following the forward model at each iteration makes these sampling methods computationally expensive and time consuming. In addition, they require some adjustments specific to each model, such as the choice of burn-in length, thinning and the number of samples to store. Harms and Roebroeck (2018) recommend to use the Adaptive Metropolis-Within-Gibbs (AMWG) algorithm for sampling dMRI models, initialized with a maximum likelihood estimator obtained from non-linear optimization, with 100 to 200 samples for burn-in and no thinning. The recommended number of samples is model-dependent. Authors recommendations can be found in their paper.
4.1.4 Bypassing the likelihood function
An alternative method was proposed to overcome the challenges associated with approximating the likelihood function and the limitations of MCMC sampling algorithms. This approach involves directly approximating the posterior distribution by using a conditional density estimator, i.e. a family of conditional probability density function approximators denoted as $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$. These approximators are parameterized by $\phi$ and accept both the parameters $\boldsymbol{\theta}$ and the observation $\boldsymbol{x}$ as input arguments. Our posterior approximation is then obtained by minimizing its average Kullback-Leibler divergence with respect to the conditional density estimator for different choices of $\boldsymbol{x}$, as per (Papamakarios and Murray, 2016):
@@ 4.1.4 Bypassing the likelihood function
If the class of conditional density estimators is sufficiently expressive, it can be demonstrated that the minimizer of Eq. (6) converges to $p(\boldsymbol{\theta}|\boldsymbol{y})$ when $N\to\infty$ (Greenberg et al., 2019). It is worth noting that the parametrization $\phi$, obtained at the end of the optimization procedure, serves as an amortized posterior for various choices of $\boldsymbol{x}$. Hence, for a particular observation $\boldsymbol{x}_{0}$, we can simply use $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x}_{0})$ as an approximation of $p(\boldsymbol{\theta}|\boldsymbol{x}_{0})$.
4.2 µGUIDE framework
The full architecture of the proposed Bayesian framework, dubbed µGUIDE, is presented in Fig. 1. The analysis codes underpinning the results presented here can be found upon publication in the Cardiff University data catalogue and on Github: https://github.com/mjallais/uGUIDE (both CPU and GPU are supported).
µGUIDE is comprised of two modules that are optimized together to minimize the Kullback–Leibler divergence between the true posterior distribution and the estimated one for every parameters of a given forward model. The ’Neural Posterior Estimator’ (NPE) module uses normalizing flows to approximate the posterior distribution, while the ’Multi-Layer Perceptron’ (MLP) module is used to reduce the data dimensionality and ensure fast and robust convergence of the NPE module. The following Sections provide more details about our implementation of each module.
4.2.1 Neural Posterior Estimator
@@
In this study, the Sequential Neural Posterior Estimation (SNPE-C) algorithm (Papamakarios and Murray, 2016; Greenberg et al., 2019) with a single round is employed to train a neural network that directly approximates the posterior distribution. Thus, sampling from the posterior can be done by sampling from the trained neural network. Neural density estimators have the advantage of providing exact density evaluations, in contrast to Variational Autoencoders (VAE) Kingma and Welling (2019) or generative adversarial networks (GAN) Goodfellow et al. (2014), which are better suited for generating synthetic data.
@@
The conditional probability density function approximators used in this project belong to a class of neural networks called normalizing flows (Papamakarios et al., 2021). These flows are invertible functions capable of transforming vectors generated from a simple base distribution (e.g. the standard multivariate Gaussian distribution) into an approximation of the true posterior distribution. An autoregressive architecture for normalizing flows is employed, implemented via the Masked Autoregressive Flow (MAF) Papamakarios et al. (2017), which is constructed by stacking five Masked Autoencoder for Distribution Estimation models (MADE) Germain et al. (2015). An explanation of how MAF and MADE work is provided in Appendix B.
To test that the predicted posteriors for a given model are not incorrect we use posterior predictive checks, which is described in more details in Appendix C.
@@
4.2.2 Handling the large dimensionality of the data with Multi-Layer Perceptron
@@
As the dimensionality of the input data $\boldsymbol{x}$ grows, the complexity of the corresponding inverse problem also increases. Accurately characterizing the posterior distributions or estimating the tissue microstructure parameters becomes more challenging. As a consequence, it is often necessary to rely on a set of low-dimensional features (or summary statistics) instead of the raw data for the inference task process (Blum et al., 2013; Fearnhead and Prangle, 2012; Papamakarios et al., 2019). These summary statistics are features that capture the essential information within the raw data, allowing to reduce the size of the input vector. Learning a set of sufficient statistics before estimating the posterior distribution makes the inference easier and offers many benefits (see e.g. the Rao-Blackwell theorem).
@@
A follow-up challenge lies in the choice of suitable summary statistics. For well-understood problems and data, it is possible to manually design these features using deterministic functions that condense the information contained in the raw signal into a set of handful summary statistics. Previous works, such as Novikov et al. (2018b) and Jallais et al. (2022), have proposed specific summary statistics for two different biophysical models. However, defining these summary statistics is difficult and often requires prior knowledge of the problem at hand. In the context of dMRI, they also rely on acquisition constraints and are model-specific.
@@
In this work, the proposed framework aims to be applicable to any forward model and be as general as possible. We therefore propose to learn the summary statistics from the high-dimensional input signals $\boldsymbol{x}$ using a neural network. This neural network is referred to as an embedding neural network. The observed signals are fed into the embedding neural network, whose outputs are then passed to the neural density estimator. The parameters of the embedding network are learned together with the parameters of the neural density estimator, leading to the extraction of optimal features that minimize the uncertainty of $p(\boldsymbol{\theta}|\boldsymbol{x})$. Here, we propose to use a MLP with three layers as a summary statistics extractor. The number of features $N_{f}$ extracted by the MLP can be either defined a priori or determined empirically during training.
4.2.3 Training µGUIDE
To train µGUIDE we need couples of input vectors $\boldsymbol{x}$ and corresponding ground-truth values for the model parameters that we want to estimate, $\boldsymbol{\theta}$. The input $\boldsymbol{x}$ can be real or simulated data (e.g. dMRI signals); or a mixture of these two. We train µGUIDE by stochastically minimizing the loss function defined in Eq. (6) using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $10^{-3}$ and a minibatch size of 128. We use 1 million simulations for each model, 5% of which are randomly selected to be used as a validation set. Training is stopped when the validation loss does not decrease for 30 consecutive epochs.
4.2.4 Quantifying the confidence in the estimates
The full posterior distribution contains a lot of useful information about a given model parameter best estimates; uncertainty; ambiguity and degeneracy. To summarize and easily visualize this information, we propose three measures that quantify the best estimates and the associated confidence levels, and a way to highlight degeneracy.
We start by checking whether a posterior distribution is degenerate, that is if the distribution presents multiple distinct parameter solutions, appearing as multiple local maxima (Fig. 2). To that aim, we fit two Gaussian distributions to the obtained posterior distributions. A voxel is considered as degenerate if the derivative of the fitted Gaussian distributions changes signs more than once (i.e. multiple local maxima), and if the two Gaussian distributions are not overlapping (the distance between the two Gaussian means is inferior to the sum of their standard deviations).
For non-degenerate posterior distributions, we extract three quantities:
@@ 4.2.4 Quantifying the confidence in the estimates
Fig. 2 presents those measures on exemplar posterior distributions.
4.3 Application of µGUIDE to biophysical modelling of dMRI data
We show exemplar applications of µGUIDE to three biophysical models of increasing complexity and degeneracy from the dMRI literature. For each model, we compare the fitting quality of the posterior distributions obtained using the MLP and manually defined summary statistics.
4.3.1 Biophysical models of dMRI signal.
Model 1: Ball&Stick (Behrens et al., 2003). This is a two-compartment model (intra-neurite and extra-neurite space) where the dMRI signal from the brain tissue is modeled as a weighted sum, with weight $f_{in}$, of signals from water diffusing inside the neurites, approximated as sticks (i.e. cylinders of zero radius) with diffusivity $D_{in}$, and water diffusing within the extra-neurite space, approximated as Gaussian diffusion in an isotropic medium with diffusivity $D_{e}$. The direction of the stick is randomly sampled on a sphere. This model has the main advantage of being non-degenerate. We define the summary statistics as the direction-averaged signal (six b-shells, see Section 4.3.3).
Model 2: Standard Model (SM) (Novikov et al., 2018a). Expanding on model 1, this model represents the dMRI signal from the brain tissue as a weighted sum of the signal from water diffusing within the neurite space, approximated as sticks with symmetric orientation dispersion following a Watson distribution and water diffusing within the extra-neurite space, modelled as anisotropic Gaussian diffusion. The microstructure parameters of this two-compartment model are the neurite signal fraction $f$, the intra-neurite diffusivity $D_{a}$, the orientation dispersion index $ODI$, and the parallel and perpendicular diffusivities within the extra-neurite space $D_{e}^{\parallel}$ and $D_{e}^{\perp}$. We use the LEMONADE (Novikov et al., 2018b) system of equations, which is based on a cumulant decomposition of the signal, to define six summary statistics.
Model 3: extended-SANDI (Palombo et al., 2020). This is a three-compartment model (intra-neurite, intra-soma and extra-cellular space) where the dMRI signal from the brain tissue is modelled as a weighted sum of the signal from water diffusing within the neurite space, approximated as sticks with symmetric orientation dispersion following a Watson distribution; water diffusing within cell bodies (namely soma), modelled as restricted diffusion in spheres; and water diffusing within the extra-cellular space, modelled as isotropic Gaussian diffusion. The parameters of interest are the neurite signal fraction $f_{n}$, the intra-neurite diffusivity $D_{n}$, the orientation dispersion index $ODI$, the extra-cellular signal fraction $f_{e}$ and isotropic diffusivity $D_{e}$, the soma signal fraction $f_{s}$, and a proxy of soma radius and diffusivity $C_{s}$, defined as (Jallais et al., 2022):
$C_{s}=\frac{2}{{D_{s}}\delta^{2}}\sum_{m=1}^{\infty}\frac{\alpha_{m}^{-4}}{\alpha_{m}^{2}{r_{s}}^{2}-2}\cdot\left(2\delta-\frac{2+e^{-\alpha_{m}^{2}{D_{s}}(\Delta-\delta)}-e^{-\alpha_{m}^{2}{D_{s}}\delta}-e^{-\alpha_{m}^{2}{D_{s}}\Delta}+e^{-\alpha_{m}^{2}{D_{s}}(\Delta+\delta)}}{\alpha_{m}^{2}{D_{s}}}\right)\leavevmode\nobreak\ ,$
with $r_{s}$ and $D_{s}$ the soma radius and diffusivity respectively, and $\alpha_{m}$ the $m$th root of $(\alpha r_{s})^{-1}J_{\frac{3}{2}}(\alpha r_{s})=J_{\frac{5}{2}}(\alpha r_{s})$, with $J_{n}(x)$ the Bessel functions of the first kind. We use the six summary statistics defined in (Jallais et al., 2022), which are based on a high and low b-value signal expansion.